OpenStudy (anonymous):

1 + 2^1 + 2^2 + 2^3 + 2^4 + 2^5 + ...2^n = 2^(n+1) -1. Prove

OpenStudy (kc_kennylau):

Geometric progression

OpenStudy (anonymous):

I was told to use induction. I used induction to n=k+1 and got stuck trying to solve.

OpenStudy (anonymous):

you want to prove \[\sum_{i=0}^{n} 2^n = 2^0 + 2^1 + 2^2 + .... + 2^n = 2^{n+1} -1\] using induction

OpenStudy (anonymous):

first make a base case n = 0 so \[2^0 = 2^{0+1} - 1 = 1 \] so its true then assume an arbitrary k where \[\sum_{i=0}^{k}2^i = 2^{k+1} - 1\] is true and prove for k+1 \[\sum_{i=0}^{k+1}2^i = 2^{(k+1)+1} - 1\] do you understand what im doing so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just want to make sure is the RHS really written as: \[2^{n+1} - 1\]

OpenStudy (anonymous):

oh nvm haha i figured it out so you move on to the inductive step first split the LHS \[2^{k+1} + \sum_{i=0}^{k}2^i = 2^{k+2} - 1\] then by inductive hypothesis: \[2^{k+1} + (2^{k+1} - 1) = 2^{k+2} - 1\] \[2^{k}2 + 2^{k}2 - 1 = 2^{k+2} - 1\] \[2(2^{k} + 2^{k}) - 1 = 2^{k+2} - 1\] \[2(2^{k+1}) - 1 = 2^{k+2} - 1\] \[2^{k+2} - 1 = 2^{k+2} - 1\] Since both sides are equal this holds for k+1 thus this holds for n+1 by induction